### v1

# Truly Assessing Fluid Intelligence of Large Language Models through Dynamic Reasoning Evaluation

Abstract
Recent advances in large language models (LLMs) have demonstrated impressive reasoning capacities that mirror human-like thinking. However, whether LLMs possess genuine fluid intelligence (i.e., the ability to reason abstractly and generalize rules in novel situations) remains an open question. Existing reasoning benchmarks either focus on domain-specific knowledge (crystallized intelligence) or lack interpretability. To address these limitations, we propose DRE-Bench, a dynamic reasoning evaluation benchmark grounded in a hierarchical cognitive framework. DRE-Bench consists of 36 abstract reasoning tasks organized across four cognitive levels, with each task featuring multiple dynamic variants that test the same underlying latent rule. This design enables fine-grained, interpretable, and reliable assessments of fluid intelligence. We evaluate a range of state-of-the-art LLMs, including both general LLMs (GPT-4o, Claude 3.7) and reasoning LLMs (o1, DeepSeek-R1, QwQ, Skywork-OR1). Experimental results reveal that although most LLMs achieve competent and robust performance in low-level cognition, they struggle with high-level cognition and exhibit limited generalization as task complexity grows. Our findings highlight the gap between current LLMs and true human-like fluid intelligence and offer a new path for systematically tracking reasoning progress in LLMs.
11footnotetext: Equal contribution22footnotetext: Corresponding author.
1 Introduction
In recent years, large language models (LLMs) openai2024gpto1 ; DeepSeekAI2025DeepSeekR1IR ; 2024claude ; openai2024gpt4o ; yang2024qwen2 have achieved remarkable success across various applications, such as intelligent chatbots zhang2023chatgpt ; ouyang2022training and code generation chen2021evaluating ; nijkamp2023codegen . Models such as OpenAI’s O1 and DeepSeek’s R1 leverage substantial test-time computation to refine their reasoning processes, explore diverse strategies, and learn from previous errors, exhibiting a degree of cognitive flexibility and self-correction that closely mirrors aspects of human-like thinking. Such advanced reasoning abilities are critical for solving complex problems in disciplines hendrycks2021math ; cobbe2021gsm8k ; lewkowycz2022minerva and code competitions chen2021codex ; li2022alphacode , and are poised to revolutionize scientific discovery wu2024tamgen ; bran2024chemcrow . As such, there is an urgent need for a rigorous and principled evaluation framework to systematically track and quantify the reasoning intelligence of cutting-edge LLMs.
Figure 1: (a) Examples of the latent rule that models are expected to understand and apply. (b) Comparison with previous benchmarks, showing that our DRE-Bench aligns with cognitive levels and controls complexity via variables. (c) Leaderboard of LLM intelligence on DRE-Bench, with accuracy on the vertical axis and stability on the horizontal axis. Higher accuracy and lower variance indicate stronger fluid intelligence.
Existing evaluation benchmarks in reasoning can be broadly categorized into two major types: crystallized intelligence cattell1963theory ; schipolowski2014nature and fluid intelligence cattell1963theory ; kent2017fluid . Crystallized intelligence refers to a model’s ability to apply accumulated knowledge to solve problems. Representative benchmarks such as AIME ye2025aime , GPQA rein2024gpqa , and SuperGPQA du2025supergpqa consist of high-difficulty discipline questions that require multi-step reasoning grounded in domain-specific knowledge. However, as models increasingly achieve expert-level performance on such knowledge-intensive tasks, it becomes increasingly important to evaluate their fluid intelligence, i.e., the ability to generalize beyond memorized content and reason in novel settings flanagan2007cattell . For example, benchmarks like ARC-AGI chollet2024arc and Raven’s Progressive Matrices raven2003raven assess this capacity through abstract reasoning tasks, i.e., infer the latent rule solely from provided input-output pairs and generalize it to predict correct outputs for novel inputs. Examples of the discovery of latent rules, such as frequency identification and category classification, are illustrated in Figure 1(a).
Although these benchmarks can measure fluid intelligence through task performance scores, they face several limitations. First, the latent rules are typically implicit and lack systematic alignment with human cognitive abilities. Most benchmarks chollet2019measure ; wu2025understanding ; yu2025stochastic present a collection of abstract reasoning tasks without explicitly aligning them with distinct dimensions or stages of human cognition primi2001complexity . As a result, it is difficult to determine how a model’s reasoning ability corresponds to specific levels of human-like intelligence. Second, each task is usually associated with a distinct latent rule, which introduces ambiguity in assessing the model’s true reasoning capability. Even if an LLM completes a task, it remains unclear whether it has genuinely understood and reasoned about the underlying rule or merely exploited superficial cues memorized during pre-training. Although recent efforts wu2025understanding ; yu2025stochastic have attempted to provide more fine-grained assessments of fluid intelligence, such as analyzing atomic operations wu2025understanding , these approaches still fall short in truly measuring fluid intelligence.
To address these challenges, we propose a Dynamic Reasoning Evaluation benchmark, DRE-Bench, designed to assess the genuine fluid intelligence of large language models (LLMs). DRE-Bench is structured around a four-level cognitive hierarchy, from simple to complex reasoning. Each level assesses 3 latent rules specified by 3 abstract reasoning tasks. Each task includes multiple dynamic variants, all designed to evaluate the same underlying latent rule. In total, DRE-Bench provides over 3,000 abstract reasoning cases. This design enables fine-grained, cognition-aligned evaluation, and allows for a robust assessment of fluid intelligence by analyzing both accuracy and variance across rule-consistent task variants.
Compared to existing benchmarks, DRE-Bench offers two key advantages as illustrated in Figure. 1(b). i) Cognition-aware task design. Unlike benchmarks such as ARC chollet2019measure and Atomic ARC wu2025understanding that present reasoning tasks without cognitive structure, DRE-Bench explicitly aligns each task with four human-like cognitive levels: attribute, spatial, sequential, and conceptual. This alignment provides good interpretability and allows mapping model behavior to specific cognitive capabilities. ii) Verifiable dynamic evaluation. To assess whether LLMs have truly grasped the underlying rules, DRE-Bench supports dynamic generation of multiple task instances by flexibly varying the latent rule-related variables. Due to the simple data format with numerical grids, all dynamic task samples and their answers can be generated deterministically via code. This ensures high reliability and correctness of evaluation data, a property that is difficult to guarantee in previous dynamic benchmarking approaches zhu2023dyval ; fan2023nphardeval ; zhu2024dynamic ; yang2024dynamic .
We conduct comprehensive experiments on our proposed DRE-Bench with various LLMs, including general LLMs (models without thinking) such as GPT-4o openai2024gpt4o and Claude-3.7 2024claude , and reasoning LLMs (models with thinking) such as OpenAI-o1 openai2024gpto1 , DeepSeek-R1 DeepSeekAI2025DeepSeekR1IR , QwQ yang2024qwen2 , Skywork-OR1 skywork-or1-2025 , etc. The takeaways of our key findings are as follows:
  • As the cognitive level of the reasoning tasks increases, model accuracy consistently declines, particularly for tasks involving physical concepts. Among them, OpenAI-o1 and DeepSeek-R1 demonstrate stronger performance and stability, while Claude 3.7 stands out in general LLMs as shown in Figure 1(c) and Section 4.2.
  • Reasoning LLMs outperform general LLMs on most abstract reasoning tasks. Moreover, as the cognitive level increases, the difference between models becomes more pronounced: differences may be minimal on lower-level tasks, but in higher-level tasks, stronger LLMs will exhibit a more obvious advantage (Section 4.2).
  • We analyzed model accuracy and stability across different complexities. We observed that with the complexity of a specific task increasing, models whose performance declines may not possess genuine fluid intelligence; only those that continue to perform well can be considered to truly master the underlying reasoning rules (Section 4.3).
  • Increasing the number of in-context training examples can slightly boost LLMs’ performance. However, adding visual information about the abstract reasoning problems has little positive impact, and sometimes even leads to a decrease in model accuracy (Section 4.4).
  • Inference time scaling plays a more important role in low-level reasoning tasks, but may be insufficient towards high-level latent rules as complexity increases (Section 4.4).
Overall, the contributions of this paper are summarized as follows. 1) We propose an abstract reasoning benchmark with a cognition hierarchy, providing a more structural and comprehensive system to analyze the LLMs’ true fluid intelligence. 2) We develop a verifiable and scalable data engine to dynamically generate abstract reasoning data with various complexities, by designing a generator and solver for each task. 3) We perform comprehensive evaluations on a variety of popular LLMs, indicating that the existing LLMs still struggle to solve the reasoning problem of high cognitive levels. Existing LLMs may not have truly internalized the underlying reasoning rules, which highlights that they remain far from achieving true fluid intelligence. Our data and code will be available at https://github.com/yangyue5114/DRE-Bench.
2 Related Work
2.1 Evaluation for Fluid Intelligence
There has been an abundance of attempts to define and measure the intelligence degree of existing large language models. Among them, the Abstraction and Reasoning Corpus(ARC)  chollet2019measure is regarded as a milestone, which defines that true intelligence should possess skill-acquisition efficiency. This concept attracted broad attention and led to many analytical studies wu2025understanding ; yu2025stochastic ; acquaviva2022communicating ; xu2023llms ; wang2023hypothesis ; wang2024speak .  wu2025understanding select some atomic abstract reasoning operations, and find that LLMs perform poorly on some atomic operations.  yu2025stochastic designed PHYSICO to evaluate whether LLMs really understand the physical phenomena they describe, by comparing language-format description and corresponding ARC format grid. However, existing abstraction reasoning benchmarks haven’t categorized tasks along cognitive dimensions, and can only provide a coarse-grained evaluation of LLMs’ reasoning ability. In addition, all these benchmarks are static, implying that they are highly susceptible to data contamination and only possess fixed complexity. Therefore, our work proposes DRE-Bench, a hierarchical cognitive dynamic benchmark on abstract reasoning. DRE-Bench can automatically generate data with varying levels of complexity, enabling comprehensive and fine-grained evaluation of LLM intelligence.
2.2 Dynamic Evaluation
Studies li2024latesteval ; li2024open ; yang2024dynamic have found that static benchmarks are highly prone to data contamination and have detected severe data contamination rates in some LLM benchmarks like wang2018glue ; wang2024mmlu . Moreover, their static nature implies a fixed level of complexity, making it difficult to adapt to evolving model capabilities. Therefore, some researchers have pioneered the exploration of dynamic evaluation on LLMs. Study (zhu2023dyval, ) proposed DyVal to dynamically generate test samples based on the graph structure to combat data contamination. Similarly, NPHardEval (fan2023nphardeval, ) generates new evaluation samples for NP-hard mathematical problems. To extend dynamic evaluation to more diverse NLP tasks, zhu2024dynamic further developed MPA, which employs LLM-based agents to transform existing problems into new ones. However, most of these dynamic evaluation methods are designed for general NLP tasks and are not applicable to more complex reasoning scenarios. More critically, the accuracy of their dynamically generated data is difficult to verify, leaving their reliability in constant doubt. In this work, we are the first to introduce a dynamic evaluation paradigm for abstract reasoning tasks. Our data generation process is code-verifiable, ensuring 100% reliability of the generated samples.
3 Method
3.1 Constructing cognition-inspired abstract reasoning framework
Studies of fluid intelligence raven2003raven ; Carpenter2018 ; primi2001complexity indicate that the complexity of a reasoning problem may be related to the types of rules applied in the inductive reasoning process. The type of rules refers to different transformations applied to elements. Among them, the rule-type hierarchy proposed by Ricardo primi2001complexity represents a relatively comprehensive cognitive framework in psychology. This cognitive hierarchy categorizes inductive rule-type as four top-down levels, and for each level they outline several corresponding types of rule variables.
According to this cognitive hierarchy, we propose our abstract reasoning framework as shown in Figure 2. For the first-tier framework, we adopt four levels, namely (1) Attribute, (2) Spatial, (3) Sequential, and (4) Conceptual. Then, for each cognitive level, we summarize a series of related rule variables related to abstract reasoning tasks. Finally, for each rule variable, we design three sets of dynamic case generators to enable fine-grained evaluation of LLMs’ corresponding cognitive reasoning capabilities. The detailed dataset table is in Supplementary Materials.
Figure 2: Specific abstract reasoning tasks across our four cognitive levels. For each task example, we visualize two pairs of input and output, corresponding to two different values of the dynamic variable. The arrows are annotated with the value range of the variable, and darker colors indicate higher complexity.
Level-1: Attribute. In the attribute level, we follow the operational dimensions identified in cognitive psychology primi2001complexity , dynamically evaluating the reasoning capabilities of LLMs along three key rule types: size, count, and shape.
Level-2: Spatial. In the spatial level, drawing on psychological literature and ARC-related data, we designed a set of classic rules that comprehensively capture the notion of spatial reasoning, namely move, rotation, and symmetry. Specifically, for the “move” rule, we design dynamic data along five directional axes: up, down, left, right, and upper-right. For each direction, we generate five sets of data where the moving distance varies from 1 to 30. This enables a fine-grained assessment of the LLM’s understanding of both moving direction and distance. Similarly, for the “rotation” rule, we design two types of rotation axes, namely around an endpoint and around the center of objects. For each rotation setting, we generate dynamic variants by changing the rotation angle, ranging from 0 to 360 degrees. For the “symmetry” rule, we design tasks based on different axes of symmetry, covering three types: horizontal, vertical, and diagonal symmetry. For each type, the number of objects to be symmetrized can vary arbitrarily.
Level-3: Sequential. For Level-3, we incorporate reasoning rules that require multi-step inference and higher-order abstract ability. Specifically, we include: category learning, which requires identifying categories based on shared attributes across varying contexts; sorting, which requires understanding order and rearranging placement; and planning, which involves goal-directed problem solving by multiple reasoning steps. To precisely control task complexity within these reasoning types, we designed corresponding rule variables: the number of categories to be distinguished, the number of elements to be sorted, and the number of planning steps required.
Level-4: Conceptual. For Level-4, we focus on scientific concepts, which require not only high-level abstract reasoning but also the application of conceptual knowledge. Inspired by fundamental branches of physics yu2025stochastic , we construct three representative concepts: gravity, reflection, and expansion. To dynamically enhance the task complexity, we increase the times we apply physics rules.
3.2 Data generation framework
After determining the cognitive level, we proceed to select the specific rule to evaluate the LLM’s performance. To enable fine-grained and dynamic assessment of the model’s cognitive abilities, we design approximately three tasks for each rule. For example, the “move” rule includes five directional tasks such as up, down, left, right, and upper-right movement; the “rotation” rule includes three tasks: rotation around an endpoint, around the center, and rotation of objects. As shown in Figure 3, for each task, we identify its underlying constraint rule and construct a set of verifiable input-output pairs using tailored generators and solvers. This approach allows us to produce dynamic abstract reasoning cases in which variables vary within well-defined ranges.
Figure 3: DRE-Bench Data Generation Pipeline: (1) Annotators identify task-specific constraints and rules, writing a generator and solver function. (2) The generator produces dynamic input data based on different variable values. (3) The solver analyzes and computes ground-truth output for each input.
Constraint. First, for a given task case, our professional data annotators identify all case-relevant constraints as Figure 3(a), including <grid boundary>, <color assignment>, <specific structure>, <object placement>, and so on. Based on these constraints, the entire reasoning rule is encapsulated from scratch into a function pair—(generator, solver)—where dynamic variables are explicitly defined as function parameters. The generator produces diverse input grids according to different configurations of the dynamic variable parameter, while the solver derives the corresponding output grid for each input, ensuring 100% correctness of the generated data pair-(input grid, output grid).
Generator. In the second step, the generator is a task-specific function manually designed. As Figure 3(b), based on the case rules, it takes the corresponding dynamic variables as input parameters and generates the corresponding input grid. A random seed is embedded in the generator, enabling the scalable and reproducible generation of an unbounded number of diverse, constraint-satisfying samples. This automatically scalable data generation process can not only produce a large number of instances, but also ensure the trustworthiness of instances, achieving precise model evaluation.
Solver. Third, the solver is a paired function corresponding to the generator. Taking the input grid as one parameter, the solver encapsulates code that gradually parses the input grid and applies the corresponding rule to solve the task, ultimately producing the correct output grid. This step is executed entirely by code, ensuring 100% correctness of the output grid as ground-truth.
4 Experiments
In this section, we evaluate state-of-the-art large language models and investigate the following research questions through experimental results: i) How do current LLMs perform in abstract reasoning across different cognitive levels? (Section 4.2); ii) As the complexity of dynamic data increases, how will the LLM’s performance change? (Section 4.3); iii) Based on the performance of different LLMs across various cognitive dimensions, to what extent has the model’s intelligence level reached? (Section 4.3); iv) Is inference time scaling, visual information, and number of training context samples, truly effective for abstract reasoning tasks? (Section 4.4).
4.1 Experimental Settings
Evaluated LLMs. For completeness, we test 11 representative LLMs varying in parameters, vision encoders, including close-sourced APIs and open-sourced LLMs. Close-sourced APIs from different companies encompass GPT-4o openai2024gpt4o , OpenAI-o1 openai2024gpto1 , Claude-3.7 2024claude and OpenAI-o3-mini openai2025o3mini . Open-sourced LLMs include DeepSeek-R1 DeepSeekAI2025DeepSeekR1IR , QwQ, Qwen2.5 yang2024qwen2 , and Skywork-OR1 skywork-or1-2025 . See Supplementary Materials for details of evaluated LLMs. To reduce randomness, all presented results of models are average results over three trials.
Evaluation Methods. In the DRE-Bench benchmark, accuracy serves as the primary evaluation metric, defined as the proportion of samples for which the model’s output grid exactly matches the ground-truth output grid. To avoid contingency, each variable contains an average of 11 samples for each of its values. All inferences are performed using the vLLM backend kwon2023efficient . To ensure fairness and consistency, we adopt the official standardized prompting template released by ARCPrize arcprize2024modelbaseline . We conducted all inference evaluations of open-source models on 8 NVIDIA A100 (80G) GPUs, while performing evaluations of APIs on CPUs.
4.2 Main Results in Four Levels
Based on the defined cognitive levels from psychology, we first evaluate model performance at each level. The main results are presented in Table 1. Overall, as the cognitive level increases, model performance exhibits a clear downward trend, which aligns with established rules in human cognitive development. Among general LLMs, Claude-3.7 consistently achieves the highest performance across all levels. Notably, it performs well even on Level 3 tasks, where many models struggle significantly. When comparing general-purpose models with reasoning-specialized models, the latter consistently outperform the former in terms of average cognitive level. Among the reasoning models, both OpenAI-o1 and DeepSeek-R1 demonstrate clear advantages. A substantial performance gap is observed between vanilla LLMs and reasoning-enhanced LLMs—for example, QwQ-32B versus Qwen2.5-32B—showing an average difference of over 20%.
Furthermore, as task difficulty increases, performance disparities among models become more pronounced, highlighting the potential of incorporating dedicated reasoning paradigms for addressing fluid intelligence problems. For Level 3B tasks, which require conceptual knowledge, all existing models fail, underscoring the current limitations of even advanced reasoning models. These findings emphasize both the inherent challenges posed by our benchmark and its flexibility in revealing model capabilities across a wide spectrum of cognitive demands.
Table 1: Model performance across four cognitive reasoning levels and corresponding tasks.
Level 1 Attribute Level 2 Spatial Level 3 Sequential Level 4 Conceptual
Model Size Count Shape Avg-1 Rotation Move Symmetry Avg-2 Category Sort Planning Avg-3 Optics Mechanics Thermal Avg-4
General LLMs
Claude-3.7 65.22 63.14 13.33 58.76 68.57 57.80 49.33 58.43 54.44 2.50 54.44 44.05 8.00 15.87 0.00 7.96
Qwen3-32B 61.79 71.05 18.33 60.05 51.43 29.20 1.33 27.66 7.69 3.75 8.89 7.14 0.00 0.00 0.00 0.00
GPT-4o 62.81 44.48 13.33 51.2 27.30 3.80 2.67 9.9 8.89 2.50 8.89 7.61 0.00 0.00 0.00 0.00
Qwen2.5-32B 44.72 28.42 6.67 35.06 5.71 0.20 0.00 1.65 4.62 1.25 7.78 4.57 0.00 0.00 0.00 0.00
Reasoning LLMs
o1 64.75 60.00 58.33 62.45 93.08 69.69 6.67 58.88 26.67 11.25 53.33 28.92 0.00 7.94 0.00 2.65
DeepSeek-R1 60.83 69.43 8.33 57.86 82.72 78.90 16.00 62.79 44.44 0.00 44.44 35.55 0.00 1.59 0.00 0.53
o1-mini 40.33 65.43 18.33 46.25 63.04 32.10 0.00 31.78 43.33 7.50 43.33 36.16 0.00 0.00 0.00 0.00
o3-mini 31.48 60.10 71.67 45.49 50.14 20.00 1.33 23.13 25.56 7.50 25.56 21.95 0.00 31.75 0.00 10.58
QwQ-32B 78.59 61.05 13.33 65.49 64.76 22.80 4.00 29.12 12.31 0.00 34.44 14.27 0.00 0.00 0.00 0.00
SkyWork-OR1-32B 59.62 68.95 13.33 57.59 64.76 15.90 4.00 25.98 9.23 0.00 36.67 12.87 0.00 0.00 0.00 0.00
4.3 Dynamic Trends across Different Cognitive Levels
Since our generator is capable of producing data with varying levels of complexity, we conduct a fine-grained evaluation to assess model performance across different cognitive levels. Figure 14 illustrates representative performance curves corresponding to each cognitive level. Additional results are provided in Supplementary Materials.
In Level-1, models exhibited relatively uniform performance patterns on the count task. Since these tasks involve basic enumeration without substantial cognitive demands, the models consistently achieved high average accuracy, and increases in complexity had minimal impact.
For Level-2, performance gaps among models became increasingly pronounced. Even as the complexity of these spatial tasks decreased, lower-performing models continued to struggle considerably. Impressively, models that excelled in overall evaluation remained robust and relatively unaffected by complexity increase, maintaining stable accuracy with minor fluctuations. This observation suggests that these models have, to some extent, acquired the capability to resolve spatial reasoning problems.
Regarding tasks in Level-3, we observe a substantial performance drop as the number of required planning steps increases. Most models can only manage the simplest scenarios, with a consistent failure point emerging when the planning depth reaches two steps. This highlights substantial limitations in the current intelligence of LLMs, suggesting that they have not genuinely mastered these planning-based rules.
Finally, at Level-4, even when the complexity of the cases is reduced, models still fail to provide correct solutions. This finding indicates that current models have only a rudimentary grasp of physical concepts and have yet to internalize even the most fundamental principles of intuitive physics.
Figure 4: Model performance curves under varying complexities in four cognitive reasoning levels.
To further illustrate the performance and stability of each model on dynamic tasks, Figure 5 presents the mean accuracy and corresponding variance across different cognitive levels. As shown in the figure, for the majority of Level-1 tasks, OpenAI-o1, DeepSeek-R1, and Claude-3.7 demonstrate strong performance and high stability. However, when the task difficulty increases to Level-2, Claude-3.7 exhibits substantial fluctuations in performance, indicating limited generalization capabilities at this level. In contrast, OpenAI-o1 and DeepSeek-R1 maintain comparable performance and stability to those observed at Level-1, highlighting the advantage of reaoning models in solving more cognitively demanding tasks. Moreover, in the third panel of Figure 5, most of the scatter points are concentrated in the lower-left region, suggesting that current models struggle to generalize effectively across the more complex and varied tasks at Level-3.
Figure 5: Scatter plots of model accuracy versus variance in cognitive reasoning levels and corresponding tasks, where points closer to the upper-left indicate higher accuracy and greater stability.
4.4 Ablation Study
Impact of the Number of In-context Learning Samples. Previous work brown2020language ; openai2023gpt4 has demonstrated the effectiveness of in-context learning in enhancing model performance across a variety of tasks. In this section, we vary the number of training examples to investigate how the quantity of in-context samples affects performance in our specific scenario. The results are shown in Figure 7. Overall, increasing the number of in-context samples helps models better capture underlying rules and improve performance. In higher levels like Level-2, Level-3 and Level-4, increasing the number of in-context training samples leads to noticeable performance improvements. However, for Level-1 tasks, increasing the number of samples yields limited improvement. This suggests that adding more in-context examples has limited impact when the model has already mastered the task or lacks the inherent capability to solve it. In both cases, the effectiveness of in-context learning is diminished.
Impact of the Auxiliary Visual Information. Previous studies legris2024h ; patterson2014human have shown that humans tend to perform better on abstract reasoning tasks when the grids are visualized, as visualization can aid in recognizing patterns and rules. Motivated by these findings, we investigate whether adding auxiliary visual information can enhance model performance. Specifically, we visualize each case and experiment with two formats: single-image, which presents all three training input-output pairs along with the test input in a single image; multi-image, which provides them as seven separate images. Table 7 presents the experimental results of GPT-4o and Claude 3.7 across cases spanning all four cognitive reasoning levels. Overall, neither visual format consistently outperforms the text-only baseline, and in some instances, accuracy even declines. These results suggest that current models struggle to derive meaningful improvements in abstract reasoning from visualized inputs.
Figure 6: Accuracy of DeepSeek-R1 on different numbers of in-context training samples.
Figure 7: Accuracy comparison between text-only(-), single-image(single), and multi-image(multi) settings, across four cognitive reasoning levels.
Model Vision L-1 L-2 L-3 L-4
GPT-4o - 88.42 2.86 5.00 0.00
Single 78.95 1.44 0.00 0.00
Multi 74.74 8.57 5.00 0.00
Claude-3.7 - 95.26 25.71 45.00 15.87
Single 96.84 17.14 31.25 15.87
Multi 97.89 17.14 35.00 12.70
Impact of the Inference Time. Recent studies DeepSeekAI2025DeepSeekR1IR ; openai2024gpto1 ; qin2024o1 ; huang2024o1 have demonstrated that inference-time scaling plays a crucial role in enhancing model performance on complex tasks. Building upon these, we take a step to examine how the model’s inference time varies with increasing task complexity. Since the internal reasoning process of O1 is inaccessible, we use the response latency to measure for the inference time. The results are presented in Figure 9. We observe that as task complexity increases, the model tends to engage in deeper reasoning. However, in tasks at Level-3 (i.e., planning), where the model can only solve problems of relatively low complexity, simply increasing inference time is insufficient to compensate for the model’s inherent limitations.
Figure 8: Changing trend in o1’s accuracy and inference time as task complexity increases.
Figure 9: Results of directional move and symmetry.
Move Symmetry
Model Up Down Left Right Horizontal Vertical
DeepSeek-R1 91.0 94.5 88.5 85.0 48 0
o1 80.0 86.5 76.5 77.0 12 8
Claude-3.7 82.0 95.0 48.0 44.0 52 36
o1-mini 15.0 34.0 53.5 57.5 0 0
Qwen3-32B 52.0 54.5 22.5 16.5 4 0
o3-mini 7.5 20.0 34.0 38.5 0 4
QwQ-32b 28.5 17.0 35.5 32.0 12 0
SkyWork-OR1-32B 5.5 4.5 31.0 37.5 12 0
GPT-4o 3.0 8.5 2.0 5.5 8 0
Qwen2.5-32B 1.0 0.0 0.0 0.0 0 0
4.5 Case Study
Analysis of Spatial Orientations. Upon closer examination of the results, we find that current models may demonstrate a distinct understanding of spatial orientation compared to humans. As shown in Table 9, the models achieve higher and more consistent accuracy in vertical (up/down) directions than in horizontal (left/right) ones in Move. Similarly, in symmetry tasks, performance is better for horizontal symmetry than for vertical symmetry. However, from the perspective of human cognition, directional distinctions are typically perceived as equivalent aflalo2008four ; ambinder2009human . These findings suggest that current LLMs may exhibit systematic divergences from human cognitive patterns in processing spatial orientation.
Analysis of Error Cases. As shown in Figure 10, we randomly select error cases from four cognitive levels where we ensure the cases are of comparable complexity. Then we visualize them alongside the corresponding ground-truth outputs for comparison. In Level-1 and Level-2, the differences between the model’s incorrect predictions and the correct answers are relatively subtle—indicating that the model roughly understands the required operation. However, in Levels-3 and Level-4, the incorrect outputs become significantly more disorganized and divergent from the ground truth, suggesting a complete failure to grasp the underlying rule. This is especially evident in Level-4, where physical concepts pose substantial challenges to the models. These observations highlight that as the cognitive level increases, the nature of model errors becomes increasingly complex and unreasonable.
Figure 10: Error cases on OpenAI-o1 across four cognitive levels. For each error case example, we visualize the input, together with the ground truth and model output.
5 Conclusion and Limitation
Conclusion In this work, we present DRE-Bench, a benchmark designed to evaluate the fluid intelligence of large language models (LLMs) through abstract reasoning tasks grounded in a structured four-level cognitive hierarchy. Unlike prior benchmarks, DRE-Bench offers interpretability and dynamic robustness by probing latent rules across multiple tasks and variants at each cognitive level. Our results show that while LLMs have advanced, fluid intelligence remains elusive. DRE-Bench offers a scalable and interpretable framework for analyzing and tracking reasoning capabilities, paving the way for future models with stronger generalizable reasoning.
Limitation. Our dynamic reasoning benchmark is fully code-generated, ensuring 100% accuracy and verifiability. While this approach guarantees data quality, it requires annotators to manually implement task-specific generators and solvers, making task expansion labor-intensive. Moreover, as discussed in Section 4.5, incorrect model responses exhibit varying degrees of deviation, suggesting that accuracy alone may be inadequate and finer-grained evaluation metrics are needed to better assess reasoning quality.
Appendix A Overview of Appendix
Due to the page limitation of the manuscript, we provide more information and details about our benchmark in the following aspects:
  • Sec. B: Braoder Impact.
  • Sec. C: Details of DRE-Bench.
    • Sec. C.1: Detailed Dataset Content and Distribution.
    • Sec. C.2: Dataset Document.
  • Sec. D: Experimental Details.
    • Sec. D.1: Details of Evaluated LLMs.
    • Sec. D.2: More Dynamic Evaluation Curves.
    • Sec. D.3: Visualization Format.
    • Sec. D.3: Table of Accuracy and Stability.
Appendix B Broader Impacts
The development of DRE-Bench represents a meaningful step toward more principled, interpretable, and verifiable evaluation of fluid intelligence in large language models. By structuring reasoning tasks within a cognitively grounded hierarchy and dynamically controlling rule-related complexity, DRE-Bench enables fine-grained assessment of models’ genuine reasoning abilities rather than surface-level pattern matching or memorization.
Moreover, the benchmark emphasizes a scalable and reproducible methodology: all task samples and solutions are generated deterministically via code, ensuring full verifiability of the evaluation process. This offers a transparent framework not only for performance comparison but also for diagnosing specific weaknesses in LLM reasoning across cognitive levels. As such, DRE-Bench can serve as a foundation for future work in designing cognitively aligned models with deeper, more structured reasoning processes.
Nonetheless, optimizing solely for DRE-Bench performance may result in artificial improvements that do not reflect true reasoning competence. Therefore, we encourage the community to treat DRE-Bench as one component within a broader suite of diverse cognitive evaluations. Future research should also explore how the cognitive insights derived from structured reasoning benchmarks can inform more human-aligned training objectives, leading to AI systems that are not only more capable, but also more interpretable, robust, and trustworthy in high-stakes environments.
Appendix C Details of DRE-Bench
C.1 Detailed Dataset Content and Distribution
To provide a more concrete overview of our dataset, we present its detailed composition and distribution in the table 2 below. This includes the specific rules, tasks, and descriptions across the four cognitive levels, along with the corresponding variables, variable ranges, and the number of data samples for each task.
Table 2: Descriptions, cognitive levels, variables, value ranges, and examples of the six atomic operations used in this paper.
Level Name Description Variable Value Range Number
Attribute Size Change the size of the whole grid or one object while maintaining the rules. size {10–30} 629
Count Change the number of grids to be counted. number {2–10} 570
Shape Change the shape of an object. shape {1–10} 450
Spatial Moving Move the object several steps towards one of {Up, Down, Left, Right, Up Right}. distance {1–30} 1500
Rotation Rotate the object around the {Endpoint, Center}. angle {0°, 360°} 108
Symmetry Perform {Vertical, Horizontal, Center} symmetry of the object. number {1, 8} 75
Sequential Categorization Classify objects based on examples, and apply the corresponding rule to each category. category {1, 5} 65
Sort Rearrange objects according to a sequential rule. order {1, 9} 240
Planning Start from an object, plan and execute a path. step {1, 9} 105
Conceptual Gravity Objects in mid-air should fall downward according to gravity. number {1, 8} 63
Reflection The light reflects upon hitting walls. number {1, 8} 100
Expansion Objects expand when heated until obstructed. number {1–8} 50
Total 3955
C.2 Dataset Document
We provide comprehensive documentation of our dataset along with its intended use cases. The dataset and accompanying resources are available at the following link: https://huggingface.co/datasets/yang5114/DRE-Bench, which includes metadata, format details, and so on.
Appendix D Experimental Details
D.1 Details of Evaluated LLMs
Table 3 lists the 11 representative LLMs examined in this study. To facilitate transparent comparison, each model is annotated along four dimensions: Model Type (General models are trained for broad-domain language generation, whereas Reasoning models have undergone additional fine-tuning or alignment specifically targeting reasoning tasks.), Param (Whenever the developer discloses the parameter count, we report it verbatim. For proprietary APIs that do not reveal their scale, the entry is marked “ –– ”.), Vision Modality, and Open-source.
Table 3: Evaluated LLMs in this study with type, specification, vision modality, and open-source status
Model Name Model Type Param Vision Modality Open-source
Claude-3.7 General Multi-modal No
Qwen3-32B General 32B Text-only Yes
GPT-4o General Multi-modal No
Qwen2.5-32B General 32B Text-only Yes
o1 Reasoning Multi-modal No
DeepSeek-R1 Reasoning 671B Text-only Yes
o1-mini Reasoning Text-only No
o3-mini Reasoning Text-only(API) No
QwQ-32B Reasoning 32B Text-only Yes
SkyWork-OR1-32B Reasoning 32B Text-only Yes
D.2 More Dynamic Evaluation Curves
Since our generator is capable of producing data with varying levels of complexity, we conduct a fine-grained evaluation to assess model performance across different cognitive levels. The four figures below illustrate performance curves of all rules corresponding to each cognitive level.
Figure 11: Model performance curves under varying complexities in level-1.
Figure 12: Model performance curves under varying complexities in level-2.
Figure 13: Model performance curves under varying complexities in level-3.
Figure 14: Model performance curves under varying complexities in level-4.
In the rules in Level-1, namely size, count, and shape, the models achieved relatively high average accuracy and stable performance since these tasks involve basic enumeration without substantial cognitive demands.
For the rotation, move, and symmetry rules in Level-2, performance gaps between models became more obvious compared to Level-1. But models still remain stable in these rules, and haven’t dropped much.
Regarding tasks in Level-3, we observe a substantial performance drop as the complexity of rules increases, whether on category reasoning, sorting, or planning.
In Level-4, although the complexity of the cases is small, models still fail to provide correct solutions and consistently present low accuracy.
D.3 Visualization Format
To provide multimodal LLMs with visual information, we designed two methods for incorporating the visual modality: one using a single image, and the other using multiple images. The figure 15 below shows some examples of single-image visual information. And multi-image means giving six input and output images of training samples and one input image of a testing sample to the LLM, respectively, and telling it what these images represent.
Figure 15: Samples of visualization format to multimodal LLMs.
D.4 Detailed table
Since plotting the accuracy and variance of all models together would make the graph unclear (or: cluttered), the Table 4 lists the specific accuracy and variance for each model to supplement the scatter plot in the main text.
Table 4: Detailed model performance across reasoning tasks (Accuracy [%] / Variance)
Model Level 1: Attribute Level 2: Spatial Level 3a: Complex Level 3b: Conceptual
Size Number Shape Rotation Move Symmetry Category Sort Planning Optics Mechanics Thermal
o1-mini 69.48/0.0133 65.43/0.0058 18.33/0.0814 63.04/0.0336 32.10/0.0215 0.00/0.0 43.33/0.0154 7.50/0.0069 43.33/0.0778 0/0.0 0.00/0.0 0/0.0
o3-mini 55.37/0.0131 60.10/0.0145 71.67/0.0381 50.14/0.0471 20.00/0.0173 1.33/0.0021 25.56/0.0183 7.50/0.0019 25.56/0.0180 0/0.0 31.75/0.0 0/0.0
gpt-4o 35.20/0.0271 44.48/0.0209 13.33/0.0156 27.30/0.0328 3.80/0.0082 2.67/0.0085 8.89/0.0354 2.50/0.0019 8.89/0.0143 0/0.0 0.00/0.0 0/0.0
Claude-3.7 50.48/0.0232 63.14/0.0037 13.33/0.0889 68.57/0.0599 57.80/0.0606 49.33/0.0853 54.44/0.0392 2.50/0.0044 54.44/0.1025 8/0.2 15.87/0.3 0/0.0
deepseek-r1 76.92/0.0074 69.43/0.0015 8.33/0.0114 82.72/0.0085 78.90/0.0159 16.00/0.0299 44.44/0.0169 0.00/0.0 44.44/0.1202 0/0.0 1.59/0.1 0/0.0
o1 80.79/0.0106 60.00/0.0063 58.33/0.0447 93.08/0.0101 69.69/0.0275 6.67/0.0064 26.67/0.0415 11.25/0.0436 53.33/0.1178 0/0.0 7.94/0.0 0/0.0
qwq-32b 78.59/0.0574 61.05/0.0190 13.33/0.0889 64.76/0.0440 22.80/0.0295 4.00/0.0192 12.31/0.0430 0.00/0.0 34.44/0.1247 0/0.0 0.00/0.0 0/0.0
skywork-32b 59.62/0.0405 68.95/0.0110 13.33/0.0456 64.76/0.0740 15.90/0.0167 4.00/0.0192 9.23/0.0340 0.00/0.0 36.67/0.0844 0/0.0 0.00/0.0 0/0.0
qwen3-32b 61.79/0.0574 71.05/0.0070 18.33/0.1347 51.43/0.0790 29.20/0.0353 1.33/0.0021 7.69/0.0580 3.75/0.0100 8.89/0.0099 0/0.0 0.00/0.0 0/0.0
qwen2.5-32b 44.72/0.1156 28.42/0.0260 6.67/0.0122 5.71/0.0270 0.20/0.0002 0.00/0.0 4.62/0.0210 1.25/0.0010 7.78/0.0062 0/0.0 0.00/0.0 0/0.0